OpenStudy (bossimbacon):

Which is 216 written as a power of a number? A. 6^3 B. 3^6 C. 108^2 D. 72^3

OpenStudy (cwrw238):

HInt: 216 = 36*6

OpenStudy (bossimbacon):

@adajiamcneal it is b right just to be sure?

OpenStudy (adajiamcneal):

yes love :)

OpenStudy (bossimbacon):

thanks<3 @adajiamcneal

OpenStudy (adajiamcneal):

:) welcome

OpenStudy (bossimbacon):

i put the question wrong ugh

OpenStudy (bossimbacon):

wait it's A isn't it? cause exponents?

OpenStudy (bossimbacon):

@adajiamcneal

OpenStudy (bossimbacon):

@Destinymasha help plz

OpenStudy (adajiamcneal):

I still think its b

OpenStudy (bossimbacon):

3^6 equals 247 am i wrong?

OpenStudy (adajiamcneal):

I searched it on google and it equals 729.....so idk

OpenStudy (bossimbacon):

and 6^3 equals 216

OpenStudy (bossimbacon):

thats why i think it's a :/

OpenStudy (adajiamcneal):

yeahhh your right I just googled it haha

OpenStudy (bossimbacon):

haha ty for clarifying it <3

OpenStudy (adajiamcneal):

anytime :)

OpenStudy (cwrw238):

thats why i said 36*6 36 = 6^2